OpenStudy (anonymous):

Which is the equation of the given line in slope-intercept form? A.y=3x+2 B.y=3x-2 C.y=-3x+2 D.y=one thirdx+2

OpenStudy (anonymous):

for some reason it wont let me put the pic

OpenStudy (anonymous):

but the points are (-1,5) (1,-1)

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

that was fast wasn't it

undeadknight26 (undeadknight26):

hmm? What was fast?

OpenStudy (anonymous):

i asked another question

OpenStudy (anonymous):

can you solve it with just the points

undeadknight26 (undeadknight26):

Of course of course...Remember what we were doing with the slope formula? We had the points? Do that for me plz?

OpenStudy (anonymous):

ok -2,6

OpenStudy (anonymous):

i think

OpenStudy (anonymous):

i think the answer is either a or b

undeadknight26 (undeadknight26):

(-1,5) (1,-1) -1 - 5/1 - (1) -6/2 = -3

undeadknight26 (undeadknight26):

Do you what the means hun?

OpenStudy (anonymous):

lol

undeadknight26 (undeadknight26):

What does it mean? Tell me tell me tell me O.O I wanna know <3

OpenStudy (anonymous):

i think it means the slope will equal 3?

OpenStudy (anonymous):

-3

undeadknight26 (undeadknight26):

AYeeeeeeeeeeeee so tel me now miss/sire...Which answer is it?

undeadknight26 (undeadknight26):

y = mx + b where m = slope! Dont eva forget!

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

hopfully im right because if i am there is still hope for me

undeadknight26 (undeadknight26):

Ew...Why would it be B? That has...That has a positive slope *eliminates A, B and D*

OpenStudy (anonymous):

ooooooohhhhhhhh im stupid duh it would be c

undeadknight26 (undeadknight26):

YESHHHHHHHHHHHHHHHH You go gurl/dood!

OpenStudy (anonymous):

imaaa bro

undeadknight26 (undeadknight26):

Go brooooooooooo Ansa dem allllllllll

OpenStudy (anonymous):

and this one would be C rightWhich equations in point-slope form are equations of the line that pass through the points (4, 5) and (−3, −1)? Choose exactly two answers that are correct. A.y-3=7over6(x-1) B.y-5=6over7(x-4) C.y+1=6over7(x+3) D.

OpenStudy (anonymous):

4,5 and -3,-1

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

last one i promise

OpenStudy (anonymous):

i hate bothering ppl

undeadknight26 (undeadknight26):

What is D?

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

y-4=7over6(x-5)

OpenStudy (anonymous):

hello

undeadknight26 (undeadknight26):

hold up bro im figuring it out...Can it be more than one?

OpenStudy (anonymous):

yes sorry

undeadknight26 (undeadknight26):

If it can be more than one its B and C...I just graphed them lol

OpenStudy (anonymous):

ok thank you

undeadknight26 (undeadknight26):

np broda <3

OpenStudy (anonymous):

i past

OpenStudy (anonymous):

yay

undeadknight26 (undeadknight26):

Thats whats OS is for ;p